OpenStudy (narissa):

Find the slope of the line passing through the points -3, 4 and 2, -1

pooja195 (pooja195):

Use the slope formula \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

pooja195 (pooja195):

Use the slope formula \[\LARGE \frac{-1-4}{2--3}=slope\] \[\LARGE \frac{-1-4}{2+3}=slope\]

OpenStudy (narissa):

i got -5/-5 and -5/5

pooja195 (pooja195):

Good divide them -5/5=?

OpenStudy (narissa):

1/-1

OpenStudy (texaschic101):

-1 is your slope

pooja195 (pooja195):

@texaschic101 can you please let them do the work? .-.

OpenStudy (texaschic101):

she already basically did

pooja195 (pooja195):

1/-1 wouldnt be considered the final answer

pooja195 (pooja195):

@narissa -5 divided by 5=?

OpenStudy (narissa):

-1

pooja195 (pooja195):

good so now we know the slope is -1

OpenStudy (narissa):

k thank you